# Construction of a Basis of Kähler differentials for generic plane branches

### v1

@@ Abstract.
As an application, we give an alternative proof for a formula of Genzmer, that provides the dimension of the generic component of the moduli of analytic classes in the equisingularity class of $\Gamma$.
Key words and phrases:
plane branch, approximate roots, moduli, Kähler differentials, flows of holomorphic vector fields
@@ 1. Introduction
The following, by no means exhaustive, is a summary of references in the literature. The moduli space $\tilde{\mathcal{C}}/\sim_{\mathrm{ana}}$ of a plane branch $\Gamma_{\circ}$ can be identified with the set of orbits of a regular action of a connected solvable algebraic group $G$ on an affine space ${\mathbb{C}}^{p}$ by a theorem of Ebey [10]. As a consequence, that moduli space can also be identified with a constructible subset of an affine space. There exists a Zariski open subset $U$ of ${\mathbb{C}}^{p}$ consisting of orbits of $G$ such that all orbits of $G$ in $U$ have the same dimension $d$ by a theorem of Rosenlicht [18]. Indeed, $U/G$ can be identified with a Zariski open subset of an algebraic variety $W$ of dimension $p-d$ (cf. [19, Chapter VI.1]) that represents the generic component of the moduli space $\tilde{\mathcal{C}}/\sim_{\mathrm{ana}}$. The generic dimension $p-d$ of the moduli was studied in the case of one Puiseux characteristic exponent by Delorme [9] and Briançon, Granger and Maisonobe [4] who gave an algorithm for its calculation. Genzmer provided a general formula for the generic dimension of the moduli space of any plane branch that depends on simple combinatorial data associated to the topological class of $\Gamma_{\circ}$ [13].
@@
Consider, for simplicity, the subset $\mathcal{C}$ of the topological class $\tilde{\mathcal{C}}$ of $\Gamma_{\circ}$ of plane branches $\Gamma$ whose tangent cone is different from $x=0$. It is obvious that every analytic class of an element of $\tilde{\mathcal{C}}$ intersects $\mathcal{C}$ and hence $\tilde{\mathcal{C}}/\sim_{\mathrm{ana}}=\mathcal{C}/\sim_{\mathrm{ana}}$. Given $\Gamma\in\mathcal{C}$, set ${\mathcal{I}}_{\Gamma}=\{\omega\in\hat{\Omega}(\mathbb{C}^{2},0):\Gamma(t)^{*}\omega\equiv 0\}$. We shall compute a basis of the ${\mathbb{C}}[[x]]$-module of Kähler differentials of $\Gamma$; more precisely, if $n$ is the multiplicity of $\Gamma_{\circ}$ at the origin, we shall find an Apery basis of $\Gamma$, i.e. germs of $1$-forms $\Omega_{1,\Gamma},\ldots,\Omega_{n,\Gamma}$ such that
@@
  • The values $\nu_{\Gamma}(\Omega_{1,\Gamma}),\ldots,\nu_{\Gamma}(\Omega_{n,\Gamma})$ define an Apery set of $\Lambda_{\Gamma}$ [2], i.e. the classes of $\nu_{\Gamma}(\Omega_{j,\Gamma})$ and $\nu_{\Gamma}(\Omega_{k,\Gamma})$ modulo $n$ are different for all $1\leq j<k\leq n$.
To begin with, define $\overline{\Omega}_{2j+1}=y^{j}dx$ and $\overline{\Omega}_{2j+2}=y^{j}dy$ for any $j\in\mathbb{Z}_{\geq 0}$. We prove that can choose the elements of the basis $\Omega_{1,\Gamma},\ldots,\Omega_{n,\Gamma}$ with sharp extra properties for any generic $\Gamma\in\mathcal{C}$:
minimality:
$\Omega_{j,\Gamma}\in\mathbb{C}^{*}\overline{\Omega}_{j}\oplus\oplus_{k=1}^{j-1}\mathbb{C}[[x]]\overline{\Omega}_{k}$ for any $1\leq j\leq n$;
monotonicity:
@@
Let us comment these properties. By the properties of an Apery basis, it follows that $\oplus_{k=1}^{j}{\mathbb{C}}[[x]]\Omega_{k,\Gamma}$ is a free $\mathbb{C}[[x]]$-module of rank $j$. Moreover, we have a filtration
@@
Let us remark that the minimality property is related to “maximal contact” properties in the spirit of Hironaka [17] for the resolution of algebraic singularities and of Aroca, Hironaka and Vicente [3] for the complex analytic case. Maximal contact varieties are essential for the resolution of singularities in characteristic 00. The $1$-forms $\Omega_{1,\Gamma},\ldots,\Omega_{n,\Gamma}$ have maximal contact in the following sense: $\nu_{\Gamma}(\Omega_{j,\Gamma})$ is the maximum of the $\Gamma$-orders of the $1$-forms in $\mathbb{C}^{*}\overline{\Omega}_{j}\oplus\oplus_{k=1}^{j-1}\mathbb{C}[[x]]\overline{\Omega}_{k}$ for any $1\leq j\leq n$ (Proposition 3.7).
@@
We divide our algorithm of construction of the Apery basis in $g+1$ levels, where $g$ is the genus of $\Gamma_{\circ}$, i.e. its number of Puiseux characteristic exponents. We define the terminal family of $1$-forms of level 00 as the set consisting of the elementary $1$-forms $dx$ and $dy$. Given a terminal family of level $l$, the initial family of level $l+1$, with $0\leq l<g$, is obtained by combining that terminal family of level $l$ with data of topological flavor. Then we describe a straightforward algorithm that provides a terminal family of level $l+1$. At the end of level $g$ we obtain an Apery basis.
@@
The algorithm of construction of an Apery basis preserves a very rich structure. For instance, the $\Gamma$-values of all the $1$-forms involved in the algorithm can be calculated in terms of combinatorial data of the equisingularity class of $\Gamma_{\circ}$ (Propositions 3.3 and 3.4). In particular, we obtain a combinatorial algorithm that provides the generic semimodule $\Lambda_{\Gamma}$ for $\Gamma\in\mathcal{C}$.
@@
In order to understand the properties of the $1$-forms that appear along the steps of the algorithm, we consider the contributions of the coefficients of the Puiseux expansion of $\Gamma$ one by one, for generic $\Gamma\in\mathcal{C}$. A similar idea was considered recently by Casas-Alvero [7]. This viewpoint leads us to define two new concepts of independent interest, namely the leading variable (or coefficient) and the $c$-sequence associated to some $\omega\in\hat{\Omega}(\mathbb{C}^{2},0)$ (Definition 3.8). They provide sufficient control to calculate the generic semigroup $\Lambda_{\Gamma}$, for $\Gamma\in\mathcal{C}$, without calculating the $1$-forms in the Apery basis or any other $1$-forms along the process.
The existence of an Apery basis with the minimality and monotonicity properties, for generic $\Gamma\in\mathcal{C}$, is proved in Theorem 3.1. Simplicity and richness are natural byproducts of our algorithm (Propositions 3.3 and 3.4). Moreover, the Apery bases given by Theorem 3.1 depend polynomially on $\Gamma$ for generic $\Gamma\in\mathcal{C}$.
Our second aim is to relate the semimodule of $\Gamma$ with analogue concepts for the strict transform $\tilde{\Gamma}$ of $\Gamma$ by the blow-up of the origin. We are forced to consider a slightly more general setting and to study the semimodule $\Lambda_{\Gamma}^{E}$ (cf. Definition 2.3) of orders of Kähler differentials on $\Gamma$ associated to $1$-forms that preserve $E$, where $E$ is a normal crossings divisor at the origin of $\mathbb{C}^{2}$. The construction is similar to the previous one, where $E=\emptyset$, and is studied in section 4. Theorem 4.1 is the analogue of Theorem 3.1 for this setting. Minimality, monotonicity, simplicity and richness are all used to provide a simple comparison of the semimodules $\Lambda_{\Gamma}^{E}$ and $\Lambda_{\tilde{\Gamma}}^{\tilde{E}}$ (Proposition 6.1), where $\tilde{E}$ is the total transform of $E$.
Let us provide an application of our construction. Consider the sequence $P_{0}=(0,0)\in\mathbb{C}^{2},P_{1},P_{2}\ldots$ of infinitely near points of a plane branch $\Gamma$ (Remark 5.3). Let $\pi_{j}$ be the blow-up of the point $P_{j-1}$. Denote by $\Gamma_{j}$ the strict transform of $\Gamma$ by $\tau_{j}:=\pi_{1}\circ\ldots\circ\pi_{j}$ where $\Gamma_{0}=\Gamma$ and $\tau_{0}=\mathrm{id}$. Fix $j\in\mathbb{Z}_{\geq 0}$. Now, let us consider actions provided by the flow of a vector field on $\Gamma_{j}$. Actions of vector fields as a tool for the study of the analytic classification problem were introduced by the first author in [11]. Consider a holomorphic vector field $X$ defined in a neighborhood of $P_{j}$ such that $E_{j}:=\tau_{j}^{-1}(0,0)$ is $X$-invariant but $\Gamma_{j}$ is not. We obtain a family of curves $\Gamma_{j}^{\epsilon}=\mathrm{exp}(\epsilon X)(\Gamma_{j})$ defined for $\epsilon\in\mathbb{C}$ close to 00. By blow-down by $\tau_{j}$, we get a family ${(\Gamma^{\epsilon})}_{\epsilon\in(\mathbb{C},0)}$ in the equisingularity class of $\Gamma$. The Puiseux expansion of $\Gamma^{\epsilon}$ is of the form
@@
where the map $(\epsilon,t)\mapsto\Gamma^{\epsilon}(t)$ is holomorphic. Let $\beta$ be the first exponent such that $a_{\beta}(\epsilon)$ is non-constant. Any $\beta$ obtained this way, by varying $X$, is called a flow-fanning exponents at index $j$, and its set is denoted by $\Theta_{j}(\Gamma)$ (Definition 5.6). The generic dimension of the moduli is equal to
$\sum_{j\geq 1}\sharp(\Theta_{j}(\Gamma)\setminus\Theta_{j-1}(\Gamma))$
@@
by Remark 5.11 and Proposition 5.4, where we assume that $\Gamma$ is generic in its equisingularity class $\mathcal{C}$ (otherwise we replace $\Gamma$ with a generic element). Such formula is obtained by introducing a correspondence between exponents of the Puiseux expansion of $\Gamma$ and divisors of the desingularization process (Proposition 5.1). We get a geometric interpretation of the generic dimension of the moduli in which $\Theta_{j}(\Gamma)\setminus\Theta_{j-1}(\Gamma)$ is the set of exponents of a Puiseux parametrization that can be changed by the action of a vector field defined in a neighborhood of $P_{j}$ but can not be changed by the action of a vector field defined in a neighborhood of any of the points $P_{0},P_{1},\ldots,P_{j-1}$.
@@
It is possible to compute $\Theta_{j}(\Gamma)\setminus\Theta_{j-1}(\Gamma)$ in terms of the semimodules $\Lambda_{\Gamma_{j-1}}^{E_{j-1}}$ and $\Lambda_{\Gamma_{j}}^{E_{j}}$ (Proposition 5.4) by applying the results on the correspondence between orders of Kähler differentials on $\Gamma$ and the tangency orders of vector fields with $\Gamma$ described in [12]. For the case where $\Gamma$ is generic in its equisingularity class, our construction of the Apery basis and its properties allow to calculate in parallel $\Lambda_{\Gamma_{j-1}}^{E_{j-1}}$ and $\Lambda_{\Gamma_{j}}^{E_{j}}$ (Proposition 6.2), to obtain $\sharp(\Theta_{j}(\Gamma)\setminus\Theta_{j-1}(\Gamma))$ for $j\in\mathbb{Z}_{\geq 1}$. This term is indeed the $j$th-term $\sigma(\nu_{P_{j-1}}(\Gamma_{j-1}\cup\tau_{j}^{-1}(0,0)))$ in Genzmer’ formula (Proposition 6.1) (cf. Definitions 6.2 and 2.1). So, we obtain a geometric interpretation of Genzmer’s formula for the generic dimension of the moduli space $\mathcal{C}/\sim_{\mathrm{ana}}$ (Theorem 6.1).
@@
We introduce the setting of the paper in section 2. The construction of the Apery basis of a generic plane branch in an equisingularity class is carried out in section 3. This is generalized for the case of Kähler differentials preserving a normal crossings divisor $E$ in section 4. Section 5 is devoted to introducing the correspondence between exponents of the Puiseux expansion and divisors of the desingularization process of a plane branch $\Gamma$ and to use such a setting to introduce the sets of flow-fanning exponents and describe their properties. Section 6 gives, as a direct application of our results, an alternative proof of Genzmer’s formula.

### v3

# Explicit Computation of The Generic Component of the Analytic Moduli of a Plane Branch

@@ Abstract.
As a first application, we give an algorithm providing the generic semimodule in an equisingularity class in terms of its multiplicity and its Puiseux characteristic exponents. As another application, we give an alternative proof for a formula of Genzmer, that provides the dimension of the moduli of analytic classes in the equisingularity class of $\Gamma$.
@@ 1. Introduction
The following, by no means exhaustive, is a summary of references in the literature. The moduli space $\tilde{\mathcal{C}}/\sim_{\mathrm{ana}}$ of a plane branch $\Gamma_{\circ}$ can be identified with the set of orbits of a regular action of a connected solvable algebraic group $G$ on an affine space ${\mathbb{C}}^{p}$ by a theorem of Ebey [10]. As a consequence, that moduli space can also be identified with a constructible subset of an affine space. There exists a Zariski open subset $U$ of ${\mathbb{C}}^{p}$ consisting of orbits of $G$ such that all orbits of $G$ in $U$ have the same dimension $d$ by a theorem of Rosenlicht [18]. Indeed, $U/G$ can be identified with a Zariski open subset of an algebraic variety $W$ of dimension pdp-d (cf. [19, Chapter VI.1]) that represents the generic component of the moduli space $\tilde{\mathcal{C}}/\sim_{\mathrm{ana}}$. The generic dimension pdp-d of the moduli was studied in the case of one Puiseux characteristic exponent by Delorme [9] and Briançon, Granger and Maisonobe [4] who gave an algorithm for its calculation. Genzmer provided a general formula for the generic dimension of the moduli space of any plane branch that depends on simple combinatorial data associated to the topological class of $\Gamma_{\circ}$ [13].
@@
Consider, for simplicity, the subset $\mathcal{C}$ of the topological class $\tilde{\mathcal{C}}$ of $\Gamma_{\circ}$ of plane branches $\Gamma$ whose tangent cone is different from $x=0$. Clearly every analytic class of an element of $\tilde{\mathcal{C}}$ intersects $\mathcal{C}$ and hence $\tilde{\mathcal{C}}/\sim_{\mathrm{ana}}=\mathcal{C}/\sim_{\mathrm{ana}}$. Given $\Gamma\in\mathcal{C}$, set ${\mathcal{I}}_{\Gamma}=\{\omega\in\hat{\Omega}(\mathbb{C}^{2},0):\Gamma(t)^{*}\omega\equiv 0\}$. We shall compute a “nice” basis of the ${\mathbb{C}}[[x]]$-module of Kähler differentials for generic $\Gamma\in\tilde{\mathcal{C}}$. Specifically:
Theorem 1.1.
There is a nice basis for $\mathcal{C}$}: If $n$ is the multiplicity of $\Gamma_{\circ}$ at the origin, there exists an open subset $U$ of $\tilde{\mathcal{C}}$ and germs of $1$-forms $\Omega_{1,\Gamma},\ldots,\Omega_{n,\Gamma}$, for $\Gamma\in U$, such that
@@
  • The values $\nu_{\Gamma}(\Omega_{1,\Gamma}),\ldots,\nu_{\Gamma}(\Omega_{n,\Gamma})$ define an Apery set of $\Lambda_{\Gamma}$ [2], i.e. the classes of $\nu_{\Gamma}(\Omega_{j,\Gamma})$ and $\nu_{\Gamma}(\Omega_{k,\Gamma})$ modulo $n$ are different for all $1\leq j<k\leq n$;
  • $\Omega_{j,\Gamma}\in\mathbb{C}^{*}\overline{\Omega}_{j}\oplus\oplus_{k=1}^{j-1}\mathbb{C}[[x]]\overline{\Omega}_{k}$ for any $1\leq j\leq n$, where $\overline{\Omega}_{2k+1}=y^{k}dx$ and $\overline{\Omega}_{2k+2}=y^{k}dy$ for any $k\in\mathbb{Z}_{\geq 0}$;
@@
Moreover, $\Omega_{j,\Gamma}$ depends polynomially on $\Gamma\in U$ and $\nu_{\Gamma}(\Omega_{1,\Gamma}),\ldots,\nu_{\Gamma}(\Omega_{n,\Gamma})$ do not depend on $\Gamma\in U$.
The first two properties determine that $\Omega_{1,\Gamma},\ldots,\Omega_{n,\Gamma}$ is an Apery basis [2] of $\Gamma$ for $\Gamma\in U$. The last two itens provide properties of minimality and monotonicity of the Apery basis respectively. Let us expand on this remark. By the properties of an Apery basis, it follows that $\oplus_{k=1}^{j}{\mathbb{C}}[[x]]\Omega_{k,\Gamma}$ is a free $\mathbb{C}[[x]]$-module of rank $j$. Moreover, we have a filtration
@@
Let us remark that the minimality property is related to “maximal contact” properties in the spirit of Hironaka [17] for the resolution of algebraic singularities and of Aroca, Hironaka and Vicente [3] for the complex analytic case. Maximal contact varieties are essential for the resolution of singularities in characteristic 0. The $1$-forms $\Omega_{1,\Gamma},\ldots,\Omega_{n,\Gamma}$ have maximal contact in the following sense: $\nu_{\Gamma}(\Omega_{j,\Gamma})$ is the maximum of the $\Gamma$-orders of the $1$-forms in $\mathbb{C}^{*}\overline{\Omega}_{j}\oplus\oplus_{k=1}^{j-1}\mathbb{C}[[x]]\overline{\Omega}_{k}$ for any $1\leq j\leq n$ (Proposition 3.7).
@@
We divide our algorithm of construction of the Apery basis in $g+1$ levels, where $g$ is the genus of $\Gamma_{\circ}$, i.e. its number of Puiseux characteristic exponents. We define the terminal family of $1$-forms of level 0 as the set consisting of the elementary $1$-forms dxdx and $dy$. Given a terminal family of level $l$, the initial family of level $l+1$, with $0\leq l<g$, is obtained by combining that terminal family of level $l$ with data of topological flavor. Then we describe a straightforward algorithm that provides a terminal family of level $l+1$. At the end of level $g$ we obtain the Apery basis provided by Theorem 1.1.
@@
The algorithm of construction of an Apery basis preserves a very rich structure. Indeed, we profit from our constructions to obtain that the $\Gamma$-values of all the $1$-forms involved in the construction of the Apery basis can be calculated in terms of $n,\beta_{1},\ldots,\beta_{g}$ (Propositions 3.3 and 3.4).
Theorem 1.2.
Algorithm 1 provides the generic semimodule $\Lambda_{\Gamma}$, for $\Gamma\in U$, in terms of $n$ and the Puiseux characteristic exponents $\beta_{1},\ldots,\beta_{g}$ of the equisingularity class $\mathcal{C}$.
@@
In order to understand the properties of the $1$-forms that appear along the steps of the algorithm, we consider the contributions of the coefficients of the Puiseux expansion of $\Gamma$ one by one, for generic $\Gamma\in\mathcal{C}$. A similar idea was considered recently by Casas-Alvero [7]. This viewpoint leads us to define two new concepts of independent interest, namely the leading variable (or coefficient) and the $c$-sequence associated to some $\omega\in\hat{\Omega}(\mathbb{C}^{2},0)$ (Definition 3.8). They provide sufficient control to calculate the generic semigroup $\Lambda_{\Gamma}$, for $\Gamma\in\mathcal{C}$, without calculating the $1$-forms in the Apery basis or any other $1$-forms along the process. Simplicity and richness are natural byproducts of our algorithm (Propositions 3.3 and 3.4).
1.1. Nice bases and blow-up
Our second aim is to relate the semimodule of $\Gamma$ with analogue concepts for the strict transform $\tilde{\Gamma}$ of $\Gamma$ by the blow-up of the origin. We are forced to consider a slightly more general setting and to study the semimodule $\Lambda_{\Gamma}^{E}$ (cf. Definition 2.3) of orders of Kähler differentials on $\Gamma$ associated to $1$-forms that preserve $E$, where $E$ is a normal crossings divisor at the origin of $\mathbb{C}^{2}$. The construction is similar to the previous one, where $E=\emptyset$, and is studied in section 4. Theorem 4.1 is the analogue of Theorem 1.1 for this setting. Note that Algorithm 1 (cf. Theorem 1.2) is generalized to this case. Minimality, monotonicity, simplicity and richness are all used to provide a simple comparison of the semimodules $\Lambda_{\Gamma}^{E}$ and $\Lambda_{\tilde{\Gamma}}^{\tilde{E}}$ (Proposition 6.1), where $\tilde{E}$ is the total transform of $E$.
As an application of our construction, consider the sequence $P_{0}=(0,0)\in\mathbb{C}^{2},P_{1},P_{2}\ldots$ of infinitely near points of a plane branch $\Gamma$ (Remark 5.3). Let $\pi_{j}$ be the blow-up of the point $P_{j-1}$. Denote by $\Gamma_{j}$ the strict transform of $\Gamma$ by $\tau_{j}:=\pi_{1}\circ\ldots\circ\pi_{j}$ where $\Gamma_{0}=\Gamma$ and $\tau_{0}=\mathrm{id}$. Fix $j\in\mathbb{Z}_{\geq 0}$. Now, let us consider actions provided by the flow of a vector field on $\Gamma_{j}$. Actions of vector fields as a tool for the study of the analytic classification problem were introduced in [11]. Consider a holomorphic vector field $X$ defined in a neighborhood of $P_{j}$ such that $E_{j}:=\tau_{j}^{-1}(0,0)$ is $X$-invariant but $\Gamma_{j}$ is not. We obtain a family of curves $\Gamma_{j}^{\epsilon}=\mathrm{exp}(\epsilon X)(\Gamma_{j})$ defined for $\epsilon\in\mathbb{C}$ close to 0. By blow-down by $\tau_{j}$, we get a family ${(\Gamma^{\epsilon})}_{\epsilon\in(\mathbb{C},0)}$ in the equisingularity class of $\Gamma$. The Puiseux expansion of $\Gamma^{\epsilon}$ is of the form
@@ 1.1. Nice bases and blow-up
where the map $(\epsilon,t)\mapsto\Gamma^{\epsilon}(t)$ is holomorphic. Let $\beta$ be the first exponent such that $a_{\beta}(\epsilon)$ is non-constant. Any $\beta$ obtained this way, by varying $X$, is called a flow-fanning exponents at index $j$, and its set is denoted by $\Theta_{j}(\Gamma)$ (Definition 5.6). We profit from this construction to reinterpret the formula for the generic dimension of the moduli of $\Gamma$ by Genzmer.
Theorem ([13]).
$\sum_{0\leq j<\tau}\sigma(\nu_{P_{j}}(\Gamma_{j})+\delta_{j}),$
where $P_{0},\ldots,P_{\tau}$ are infinitely near points of $\Gamma$, $\tau$ is the first index such that $\Gamma_{\tau}$ is smooth, and $\delta_{i}$ is the number of irreducible components of $E_{i}$ passing through $P_{i}$ (cf. Definition 6.2).
Theorem 1.3.
$\sum_{j\geq 1}\sharp(\Theta_{j}(\Gamma)\setminus\Theta_{j-1}(\Gamma)),$
and indeed $\sharp(\Theta_{j+1}(\Gamma)\setminus\Theta_{j}(\Gamma))=\sigma(\nu_{P_{j}}(\Gamma_{j})+\delta_{j})$ for any $j\geq 0$, where we assume that $\Gamma$ is generic in its equisingularity class $\mathcal{C}$ (otherwise we replace $\Gamma$ with a generic element).
@@
We get a geometric interpretation of the generic dimension of the moduli space $\mathcal{C}/\sim_{\mathrm{ana}}$ in which $\Theta_{j}(\Gamma)\setminus\Theta_{j-1}(\Gamma)$ is the set of exponents of a Puiseux parametrization that can be changed by the action of a vector field defined in a neighborhood of $P_{j}$ but can not be changed by the action of a vector field defined in a neighborhood of any of the points $P_{0},P_{1},\ldots,P_{j-1}$. Thus, in particular we obtain a geometric interpretation of Genzmer’s formula. It is obtained by introducing a correspondence between exponents of the Puiseux expansion of $\Gamma$ and divisors of the desingularization process (Proposition 5.1).
@@
It is possible to compute $\Theta_{j}(\Gamma)\setminus\Theta_{j-1}(\Gamma)$ in terms of the semimodules $\Lambda_{\Gamma_{j-1}}^{E_{j-1}}$ and $\Lambda_{\Gamma_{j}}^{E_{j}}$ (Proposition 5.4) by applying the results on the correspondence between orders of Kähler differentials on $\Gamma$ and the tangency orders of vector fields with $\Gamma$ described in [12]. For the case where $\Gamma$ is generic in its equisingularity class, our construction of the Apery basis and its properties allow to calculate in parallel $\Lambda_{\Gamma_{j-1}}^{E_{j-1}}$ and $\Lambda_{\Gamma_{j}}^{E_{j}}$ (Proposition 6.2), to obtain $\sharp(\Theta_{j}(\Gamma)\setminus\Theta_{j-1}(\Gamma))$ for $j\in\mathbb{Z}_{\geq 1}$ and then show that it is the $j$th-term $\sigma(\nu_{P_{j-1}}(\Gamma_{j-1}\cup\tau_{j}^{-1}(0,0)))$ in Genzmer’ formula (Proposition 6.1) (cf. Definitions 6.2 and 2.1).
@@
We introduce the setting of the paper in section 2. The construction of the Apery basis of a generic plane branch in an equisingularity class is carried out in section 3. This is generalized for the case of Kähler differentials preserving a normal crossings divisor $E$ in section 4, where we also provide Algorithm 1 that gives the generic semimodule in an equisingularity class. Section 5 is devoted to introducing the correspondence between exponents of the Puiseux expansion and divisors of the desingularization process of a plane branch $\Gamma$ and to use such a setting to introduce the sets of flow-fanning exponents and describe their properties. Section 6 gives, as a direct application of our results, an alternative proof of Genzmer’s formula.